(1 point) Take the Laplace transform of the following initial value problem and solve for Y(8) = L{y(t)}; ſ1, 0<t<1 y – 6y
Answer #1

y cul 974 2 0 - 64 - 274 - no 1st 1 9(0)=0 y 1000 Taking Laplace transform on bothe side 224) – 62294 – 27 L3g = sestadt124) I s (57-65-27) S(57-65-27) 1 y(s) = S (5²65-27) 5(5²65-27 Mow taking invese da place 4 rll & .) - 65-97 2.27s toiset tar, -3t of Jy (t) = - I ult) + its e + Toot - 4LE-D ( te glt-1) -377-7 et Rate you et Plose ! Alor comments have problem.